Can you please help me set this problem up:

Anita’s desk is a rectangle that is twice as long as it is wide. If the area of the desktop is 18 ft2, find the length and the width of the desk.

Answer by richwmiller(17219)

L*W=18
L=2W
2w*w=18
w^2=9
w=3
l=6
yes you guessed it 6*3=18
Length is 6 and width is 3